Session: Oct 2007 (session 3)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (A. Polatidis at Tue Oct 30 11:42:11 2007) Minor failures :
Lost the first scan. For some time during the first 30min of the schedule, communication with the frontend computer was lost and the low amplitude noise diode was injected with an 80 Hz tone. This did not affect the fringes but will result in reduction in sensitivity for that time. Lost 30sec a few times during scans to make a system temperature measurement. The external filters were switched on without any problems.

Hartebeesthoek (Jonathan Quick at Tue Oct 30 11:48:32 2007) Success :
No known problems - perfect weather

Medicina (Zacchiroli at Tue Oct 30 11:50:46 2007) Success :
During all the experiment, Tsys value on IF1 and IF2 very hight. Bad weather conditions, heavy rain during all the experiment.

Torun (K. Borkowski at Tue Oct 30 11:59:53 2007) Minor failures :
There were problems with Mk5 disk2file so no data could be ftp-ed.

Onsala (Michael Lindqvist at Wed Oct 31 07:52:49 2007) Success :
No known problems.

Westerbork (Tony Foley at Wed Oct 31 08:25:31 2007) Minor failures :
Single dish was used. Please change the coordinates in the vexfile accordingly. problems with disk2file, so it had to be repeated after the end of N07M3. Also started about 20seconds late.

Noto (P. Cassaro at Wed Oct 31 10:51:43 2007) Minor failures :
Failed data ftp due to lacking of room in the local disks.

Jodrell_Bank (Alastair Gunn at Fri Nov 2 11:51:54 2007) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Fri Nov 2 11:52:37 2007) Failure :
Formatter lost sync prior to this experiment. Reset at 09:39 UT. Polarisations were swapped, corrected after n07m3. No other known problems.